A (g) + B(g) = C(g) + D(g) Above equilibrium is established by taking A and B in a closed container. Initial concentration of A is twice the initial concentration of B. At equilibrium concentration of B and C are equal. Then find the equilibrium constant for the reaction, C(g) + D(g) = A(g) + B(g)

Grade:11

1 Answers

jitender lakhanpal
10 years ago

hi chase PFA for solution

                      A (g) + B(g) = C(g) + D(g)

 

Initial             2a         a           0        0

Change          -x         -x         +x      +x

Equilibrium   2a-x     a-x        x       x

 

Given: At equilibrium Concentration of B=C

Thus, a-x = x

Or, 2x = a

Now:

C(g) + D(g) = A(g) + B(g)

 

Thus, K = [A][B] / [C][D]

Or, K = (2a-x)(a-x) / (x)(x)

Or, K = (4x-x)(2x-x) / x^2

Thus, K = 3x*x / x^2 = 3

Thus, Equilibrium constant of the reaction is K=3